7x-2y=-2 3x-3y=7
Solve the following system of linear equation by elimination

Answer by checkley71(8403)

7x-2y=-2 or -2y=-7x-2 or y=-7x/-2-2/-2 or y=7x/2+1
3x-3(7x/2+1)=7 or 3x-21x/2-3=7 or (6x-21x)/2=7+3 or (6x-21x)/2=10
now cross multiply 6x-21x=2*10 or -15x=-20 x=-20/-15 or x=4/3.
7*4/3-2y=-2 or 28/3-2y=-2 or -2y=-2-28/3 -2y=(-6-28)/3 or -2y=-34/3
or y=-34/3*1/-2 or y=34/6 or y=17/3